OpenStudy (anonymous):

4 1/9 + 7 4/15 write as a mixed numeral

myininaya (myininaya):

\[4+\frac{1}{9}+7+\frac{4}{15}=4+7 +\frac{1}{9}+\frac{4}{15}=11+\frac{1}{9}+\frac{4}{15}\] try combining the fractions and finding the sum of 1/9 and 4/15: do \[\frac{1}{9}+\frac{4}{15}\] then we will come back to adding 11

OpenStudy (turingtest):

do you understand so far???

OpenStudy (anonymous):

yeah sorta

OpenStudy (anonymous):

is the answer 11 19/45

myininaya (myininaya):

\[\frac{(1)15+4(9)}{9(15)}=\frac{15+36}{135}=\frac{51}{135}=\frac{17}{45}\] so i got 11 \frac{17}{45}\]

myininaya (myininaya):

\[\ 11 \frac{17}{45} \]